OpenStudy (ujjwal):

A 2 kg block collides with a horizontal spring of force constant 4 N/m. The spring is compressed by 8 cm from the rest position. If the coefficient of friction is 0.5, then the velocity of block with which it collide with spring is. a)0.7 m/s b)0.6 m/s c)0.9 m/s d)1.2 m/s

OpenStudy (vincent-lyon.fr):

Use energy.

OpenStudy (ujjwal):

Where am i supposed to use the co-efficient of friction? @Vincent-Lyon.Fr

OpenStudy (vincent-lyon.fr):

You will need to work out the work done by friction. F = µN = µmg and then W = - F Δl, where Δl is compression of 8 cm.

OpenStudy (ujjwal):

should it be like this??\[\frac{1}{2}Kx^2 +\mu mg\times x=\frac{1}{2}mv^2\]It gives me an answer 0.9 m/s, which is one of the options. But i don't know the correct answer.. Is 0.9m/s correct?

OpenStudy (vincent-lyon.fr):

Yep! g=10N/kg yields v=0.9m/s

OpenStudy (ujjwal):

Thanks a lot @Vincent-Lyon.Fr ..